NetCall - Release 2

Poll-handler for Fido-Net-Systems Version 2.0c

(Last change: 02.07.95)

by Alexander Stirmlinger

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994-1995 Alexander Stirmlinger

NetCall is shareware. All rights of the program, files and documents belong completely to the author. The archive and its files (program, documents and the keyfile) must not be changed in any way. No files may be added or deleted to/from the archive.

You may charge a fee to recover distribution costs. The fee for diskette distribution may not be more than the cost to obtain a public domain diskette from Fred Fish.

1 General informations

1.1 Features of NetCall

• Simplifies the treatment of file-requests and crash-mails (pending jobs are displayed in a listview-gadget).

• Enables automatic polling on all lines of the destination system (if one line is BUSY, NetCall will automatically dial up the next line).

• Includes #?.REQ-file editor (’request file’) to create or edit .REQ-files in the OUTBOUNDdirectory.

• Searches (even when your mailer polls already) for other lines of the destination system in the nodelist (crashing or frequesting).

• Supports/simplifies multi-bossing.

• Simplifies polling on a line of your boss, which is not listed in the nodelist.

• Hides your own password when crashing or frequesting.

• Supports TrapDoor in resident or in shell-mode.

• Supports automatical operation (without user-interaction).

• User-defineable toolbox.

• ARexx-interface.

• Font-sensitive GUI.

• Opens public-screen if desired.

• User-defineable menus (user-commands).

• Extensive ways to configure.

• Many shell-arguments.

• Localized (OS2.1 and better).

• Iconify-function (including QUIET-mode).

1.2 Copyright

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994-1995 Alexander Stirmlinger

NetCall is shareware. All rights of the program, files and documents belong completely to the author. The archive and its files (program, documents and the keyfile) must not be changed in any way. No files may be added or deleted to/from the archive.

The keyfile is personalized to its owner and must not be changed, saled, hired or lended.

1.3 Warranty

NetCall is provided "AS-IS". I can’t guarantee the expected work of the program.

Though, I’m always anxious to remove all bugs and misbehaviors.

Reports, hints and suggestions are always welcome. (See also Chapter 6 [Author info], Page 19.)

  THERE IS NO WARRANTY FOR THE PROGRAMS, TO THE EXTENT PERMITTED
BY APPLICABLE LAW.  EXCEPT WHEN OTHERWISE STATED IN WRITING THE
COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAMS "AS
IS" WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED,
INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF
MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE.  THE
ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAMS IS
WITH YOU.  SHOULD THE PROGRAMS PROVE DEFECTIVE, YOU ASSUME THE
COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.
  IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN
WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY
REDISTRIBUTE THE PROGRAMS AS PERMITTED ABOVE, BE LIABLE TO YOU
FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR
CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE
THE PROGRAMS (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA
BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD
PARTIES OR A FAILURE OF THE PROGRAMS TO OPERATE WITH ANY OTHER
PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED
OF THE POSSIBILITY OF SUCH DAMAGES.

1.4 Restrictions of the unregistered version

• Nerving requester

• Nerving wait-function (appears randomly)

2 Installation

2.1 Assumptions to the hardware

NetCall does not make any assumptions to the hardware. NetCall runs on every Amiga with at least 512 KB RAM.

2.2 Assumptions to the software

NetCall runs on Amigas equipped with AmigaOS 2.04 or later. AmigaOS 3.x is, of course, supported, too.

Additionally, you need TrapDoor^{1} (not included in the NetCall-archive) as a FidoNet-mailer. Other mailers may only be used, if they use the same shell-arguments and a compatible logfile (NetCall needs to scan the TrapDoor-logfile).

2.3 Installer

The installation is easily done by Commodore-Amiga’s Installer. Just double-click on one of the install-icons (english or german language). All needed files will be copied to your hard-disk.

After having installed NetCall, you’ll have to configure NetCall to your wishes and to your FidoNet-software. Please read the configuration-chapter of this document for further information. See also Chapter 3 [Configuration], Page 7.

Installer and Installer project icon (c) Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved. Reproduced and distributed under license from Commodore.

INSTALLER SOFTWARE IS PROVIDED "AS-IS" AND SUBJECT TO CHANGE; NO WAR-RANTIES ARE MADE. ALL USE IS AT YOUR OWN RISK. NO LIABILITY OR RESPONSI-BILITY IS ASSUMED.

1 TrapDoor is copyright \begin{matrix}c \\ ◯ \end{matrix} Maximilian Hantsch and Martin Laubach

2.4 Execution from a shell

The following arguments may be passed if NetCall is executed from a shell:

POLLID,CALL/K,CONFIG/K,OUTBOUND/K,PUBSCREENNAME/K,NOREQ/S,
NOPOLL/S,NOCRASH/S,NOFREQ/S,ALLACTIVE/S,OPENBIG/S,TERMREQS/S,
TELL/S

POLLID PollID(s)

This is the ID from your bosses, where you want NetCall to poll. The IDs given here must correspond with the BOSS-entries in your ‘NetCall.cfg’ (case-sensitive). You may pass a valid AmigaDOS-pattern as an argument here (‘POLLID (BBS1|BBS2)’ would cause NetCall to poll at node BBS1 and BBS2).

CALL Node/Phone Node- or phone-number where you want NetCall to poll

(example: ‘CALL 2:246/1045’).

CONFIG Config-file Use the given configuration-file rather than the normal configuration.

(example: ‘CONFIG Work:NetCall/Config2.cfg’).

OUTBOUND Outbound-path Scan this OUTBOUND-directory rather than the directory specified in the configurationfile.

PUBSCREENNAME public-screen

Force to open NetCall’s and the mailer’s windows on this public-screen.

NOREQ

Use stdIO for error-output rather than a system-requester.

NOPOLL

Do not pick up entries where NetCall would just poll (in other words: exclude your bosses).

NOCRASH

Exclude entries where NetCall would crash mails to.

NOFREQ

Exclude entries where NetCall would file-request from.

ALLACTIVE

Mark all entries as active immediately after NetCall gets started. NetCall will start to poll immediately.

OPENBIG

Always open NetCall’s big window (two list-view-gadgets, one text-gadget).

TERMREQS

Forces NetCall to terminate requesters (even error-requesters!) automatically. This option allows automatical operation without user-interaction.

TELL

If you are running TrapDoor resident in the background, NetCall may control Trap-Door via the TrapTell-command. To enable this option, specify TELL as a shellargument and define the TrapTell-command in ‘NetCall.cfg’. See also Chapter 3 [Configuration of NetCall], Page 7.

2.5 Execution from workbench

NetCall can easily be started by double-clicking on it’s icon.

The following arguments may be passed as tool-types in the program icon:

ARGS = <shell-args> All arguments, which can be passed as shell-arguments, can be passed in the ARGStool-type, too. See also Section 2.4 [Execution from a shell], Page 4.

TOOLPRI = <priority (-127 to 128)> Taskpriority of NetCall.

6 NetCall 2.0c

3 Configuration

3.1 The configuration-file NetCall.cfg

NetCall will look for its configuration-file in the following directories:

1. Mail: (if assigned)

2. The current directory

3. The directory, where the environment-variable NETCALLCONFIG points to

4. Shell-argument CONFIG (See also Section 2.4 [Execution from a shell], Page 4.)

The configuration-file of NetCall may/must include the following parameters:

OUTBOUND "Path"

OUTBOUND-directory of your mailer (example: ‘Mail:Outbound’).

NODELIST "Path"

Directory where your nodelist can be found (example: ‘Mail:Nodelist’). If you don’t have a nodelist, delete this line from your configuration-file or mark it out as a comment.

MAILER "Path/Mailer"

TrapDoor-mailer (example: ‘Mail:Bin/TrapDoor’).

TRAPTELL "Path/TrapTell" If you want NetCall to start TrapDoor not as a shell-command, but to control Trap-Door via then TrapTell-command, you must specify the TrapTell-command here. To enable the tell-function, you have to specify TELL as a shell-argument to NetCall. See also Section 2.4 [Execution from shell], Page 4.

ARGAPPEND "Mailer shell-arguments" Arguments specified here, will be passed to the mailer, directly. To work probably NetCall and not TrapDoor should be responsible for the retries. This can either be achieved by writing ‘RETRIES 0’ in your ‘TrapDoor.cfg’ or passing ‘RETRIES 0’ here as a shell-argument to the mailer. See also Section 3.2 [Configuration of TrapDoor], Page 12. (example: ‘RETRIES 0’).

ARGAPPENDICONIFIED "Mailer shell-Arguments" Additional shell-arguments passed to the mailer, when NetCall operates in the iconified mode (example: ‘QUIET’ would cause TrapDoor to open no windows, when NetCall is iconified).

TDLOGFILE "Path/TrapDoor-LogFile"

LogFile of TrapDoor (example: ‘Mail:Logs/TrapDoor.log’).

NETCALLSTART "Shell-command"

This shell-command gets executed at NetCall’s startup.

NETCALLEND "Shell-command"

This shell-command gets executed at NetCall’s shutdown.

AFTERSESSION "Shell-command"

This shell-command gets executed after each poll.

HELPCOMMAND "Shell-command"

This shell-command gets executed if you press the <HELP>-key when the NetCallwindow is active.

(‘SYS:Utilities/AmigaGuide HELP:english/NetCall.guide PUBSCREEN NETCALL’).
STARTUPDELAY seconds

Time that should delay before NetCall starts the poll-process. This delay allows you to deselect unwanted list-entries. This option is only active when ALLACTIVE is given as a shell-argument. (‘STARTUPDELAY 0’ means no delay).

REDIALDELAY seconds

Time that should delay if all lines are BUSY at the end of the poll-list.

REDIALDELAYENTRY seconds

Time that should delay after each entry in the poll-list. This option is useful for modems with redial-restrictions.

MAXRETRIES retries

After having reached this amount of retries, NetCall will automatically deactivate all active list-entries. (Use ‘MAXRETRIES 0’, if you don’t want automatical deactivation).

MAXRETRIESCOMMAND "Shell-command"

This AmigaDOS-command gets executed, if the amount of MAXRETRIES is expired.

(example: ‘rexx:rx 'address NETCALL QUIT'’)

(This example sends the ARexx-command QUIT to NetCall. This will quit NetCall.) If MAXRETRIESCOMMAND is not specified, NetCall will deselect all entries in the poll-list.

MULTILINE

This keyword will cause NetCall to search for other nodes of your destination system(s) in the nodelist. The search will be performed by a second task, so no delay of the polling-task should be noticeable. Use NOMULTILINE if you don’t want this option.

AUTOQUIT

NetCall will be automatically terminated as soon as all jobs are processed. Use NO-AUTOQUIT if you don’t want this option.

MAILERABORT (ABORTNETCALL|CONTINUE|REQUEST)

If you quit your mailer (user-break, <ESC>), NetCall may:

• be quitted, too (MAILERABORT ABORTNETCALL)

• poll at the next node (MAILERABORT CONTINUE)

• open a requester (MAILERABORT REQUEST)

SAVELASTPOLLDATE

If this option is enabled, NetCall will display the date of the last successful call to each of your bosses in the listview-gadget. NetCall will store the needed data in a file called ‘NetCall.polldates’, which can be located in your ‘Mail:’- or in the current directory at NetCall’s startup. (NOSAVELASTPOLLDATE if you don’t want this option).

OPENPUBSCREEN "Screenname" Width Height Depth ScreenModeID Force NetCall to open its own public-screen. All values must be given in decimal format. To specifiy the ScreenModeID, you may use the ScreenModeViewer-Utility

(example: ‘NETCALL 640 256 2 167936’).
PUBSCREEN "Screenname"

Open NetCall on this public-screen. If NetCall can’t find this screen to open on, NetCall will open its window on your workbench. If you wish NetCall to open its own public-screen (OPENPUBSCREEN) you must specifiy the same name at OPEN-PUBSCREEN and PUBSCREEN (example: ‘NETCALL’):

SCREENTOFRONT

If SCREENTOFRONT is specified, NetCall will automatically bring the screen where the NetCall-window is opened to the front. (NOSCREENTOFRONT if not desired).

BACKDROPWINDOW

NetCall can optionally open an invinsible backdrop-window on its own public-screen. If you enable this option, you are able to select NetCall’s menu even if you clicked outside the NetCall main-window. This option becomes only active, if NetCall opens its own public-screen. Opening a backdrop-window does not need any memory, but it could slow down window-refreshing (esp. if many windows are open). Use NOBACK-DROPWINDOW if you don’t want a backdrop-window.

PUBSCREENMAILER "Screenname"

You can optionally force the mailer to open its windows on this public-screen (example:

‘Workbench’).

WINDOW X/Y/Width/Height Window-position/-size (big NetCall-window, incl. list-view-gadgets and text-gadget)

(example: ‘30/30/400/180’).

WINDOWZOOM X/Y/Width/Height Window-position/-size of the zoomed (little) NetCall-window. If you specifiy AUTO rather than a decimal value as the height, NetCall will automatically adapt its windowheight to your actual font-size (example: ‘30/30/400/AUTO’).

WINDOWFILEREQ X/Y/Width/Height Position/size of the REQ-file editor window.

(example: ‘50/40/190/240’).
WINDOWFILEREQZOOM X/Y/Width/Height

Position/size of the REQ-file editor window in zoomed state.

(example: ‘50/40/200/100’).

WINDOWTOOLBOX X/Y/Width/Height Position/size of the toolbox. NetCall will ignore the Width value and adapt the window-width to the actual screenfont and to the gadgets. (example: ‘50/40/0/100’).

WINDOWACTIVATE

Activate the NetCall-window when opened (use NOWINDOWACTIVATE if you don’t want to activate).

WINDOWFONTNAME "Fontname"

The Font that should be used for the list-view-gadgets (example: ‘topaz.font’).

WINDOWFONTSIZE Height

Font-size used for the list-view-gadgets (example: ‘8’).

LVITEMHEIGHT Pixel

Space between the list-view-entries (given in pixel) (example: ‘1’).

OPENBIG

If you specify OPENBIG, NetCall will always open its big window, no care if there are file-requests or crash-mails in your OUTBOUND-directory. Use NOOPENBIG if you don’t want this option.

SEARCHOPTS 5/6

Search-options (for MULTILINE). Leave it be...

SEARCHPATTERN "Pattern"

Pick up all nodes with this fido-flags to the ToDo-list. A valid AmigaDOS pattern can be given here. If you specify ‘~(#?ISDN#?)’ NetCall will pick up all fido-nodes without ISDN. If you don’t specify MULTILINE, this keyword will have no effect.

SUPRESSREQ "TrapDoor error" NetCall may supress/terminate some of its error-requesters automatically. You may specify any amount of SUPRESSREQ-commands here. (Note the "#?...#?"). (exam-

ple: ‘"#?NO CARRIER#?"’).

COMMAND "Name of the menu-entry" "Shell-command" ASYNC GADGET GAP With this keyword you are abled to define your own menu-entries in the NetCallmenustrip. You may specify any number of COMMAND-keywords here. If you add ASYNC, NetCall will launch the command as an seperate process. (example: ‘Test1

SYS:Prefs/Time’).

If you want this entry appear in the toolbox, you have to append GADGET. You can insert a gap after the gadget if you specify GAP. (example: ‘Test1 SYS:Prefs/Time

GADGET GAP’).

If you want to send ARexx-commands to NetCall itself, you have to use the ASYNCflag. Otherwise this will cause a deadlock of NetCall. (NetCall would wait until the ARexx-command is finished, but the handling of the command can’t start

while NetCall is waiting). (example: ‘"Quit NetCall" "rexx:rx 'address NETCALL
QUIT'" ASYNC GADGET’).

BOSS ID(*) Node/Phone MyAddr Password Retries BeforeSession AfterSession With the BOSS-keyword you can define all lines of your own fido-bosses. You may specify any number of BOSS-keywords. By using several different IDs you can configure your system for multi-bossing. Entries with a * behind the ID will be displayed in the NetCall-window (list-view-gadget). You may use different amounts of retries for each line of your boss (optional). If the line is BUSY, NetCall will automatically retry this line (if retries is specified). You can specify phone-numbers rather than node-numbers if your boss has unlisted lines. BeforeSession and AfterSession can be scripts, which will be executed before and after the mailer-call to the corresponding BOSS (optional). Example:

      BOSS  SO*   2:246/2248   2:246/2248.20  test  3  Export  Import
      BOSS  SO    2:246/2249   2:246/2248.20  test
      BOSS  CDS*  2:246/1045   2:246/1045.11  test  0  Export  Import
      BOSS  CDS   2:246/1046   2:246/1045.11  test
      BOSS  CDS   2:246/1047   2:246/1045.11  test  3
      BOSS  CDS   07131167890  2:246/1045.11  test

CALL ID(*) Node/Phone MyAddr Password The CALL-directive can be used to add fido-nodes to NetCall’s ToDo-list, which are no personal bosses. This might be useful if you call a system regularly, to receive data put on hold for you. The password is optional. Example:

      CALL  ID1*  2:246/1234   2:246/1045.11  test
      CALL  ID2*  2:246/5678   2:246/1045.11

ADDNODE NODE ADDNODE...

With the ADDNODE-variable you can specify additional node- or phone-numbers of nodes which are currently in NetCall’s list. This might be useful to call lines which are unlisted in the nodelist. Please notice that this additional nodes/phonenumbers are not displayed in the upper list-view-gadget. They are only displayed when their NODE-entry is selected. Example:

             ADDNODE 2:246/1045 07131/123456 07131/987654
             ADDNODE 2:246/2248 07141/123456

3.2 The configuration of TrapDoor

You should set the amount of RETRIES in your ‘TrapDoor.cfg’ to 0 (‘RETRIES 0’), because NetCall will do this from now on.

NetCall will call each line of the desired destination-system, if the other lines are BUSY.

3.3 The usage of NetCall with Spot

To use NetCall with Spot^{1} as optimized as possible, you should configure Spot as follows: Execute NetCall only from your first boss-entry in the Spot-configuration (Settings -> Sy-

stem...). Enable the asynchronous-button.

   NetCall >NIL: OPENBIG

All other NetCall-entries of Spot should be cleared. If you press the NetCall-button of Spot, NetCall will be started. If you have multiple bosses, Spot will no longer ask you where to poll. You can choose this

directly in the NetCall-window. 1 Spot is copyright \begin{matrix}c \\ ◯ \end{matrix} Nico François

3.4 The usage of NetCall with UMS

If you want to use NetCall with UMS^{2} you don’t have to change anything in your ‘ums.config’.

NetCall can easily be started by double-clicking its icon on the workbench or executing it in a shell.

3.5 The usage of NetCall with MailManager

Just like at the usage of NetCall with UMS, you have to change nothing in your configuration of MailManager^{3} if you want to use NetCall with MailManager. See also Section 3.4 [The usage of NetCall with UMS], Page 13.

NetCall can easily be started by double-clicking its icon on the workbench or executing it in a shell.

2 UMS is copyright \begin{matrix}c \\ ◯ \end{matrix} Martin Horneffer 3 MailManager is copyright \begin{matrix}c \\ ◯ \end{matrix} Pino Aliberti 14 NetCall 2.0c

4 ARexx-interface

The ARexx-interface of NetCall can be reached via the public port NETCALL.

The interface knows the following commands:

ABORTDELAY

abort the actual delay-loop (if there is one).

ABORTMAILER

cancels the mailer, if he is currently active. If NetCall is in a delay-loop, this loop will be canceled, too.

BEEP flashes the screen of NetCall.

DESELECTALL

deselects all the active entries in the list-view-gadget. The poll-process of the active entry will not be stopped. (like the menu-item ’deselect all’).

HELP call the help-function of NetCall (just like the ’help’ menu-item).

ICONIFY iconifies NetCall.
ISICONIFIED

returns warn (5) if the user-interface of NetCall is currently open or ok (0) if NetCall is iconified.

NETCALL2FRONT

brings the screen of NetCall to the front.

NODEMANUAL NODE,ACTIVE/S Adds a node to NetCall’s poll-list. The poll-process will be started, if you specify ACTIVE.

Example: ‘NODEMANUAL 2:246/1045 ACTIVE’

When no parameter ist specified, NetCall will open the ’add node manual’-window.

QUIT terminates NetCall. The poll-process will be terminated, too.

REQUESTFILE

open the file-requester (just like the ’request file...’ menu-item). This command has no effect when NetCall is in ICONIFY-mode.

SELECTALL

selects all the entries in the list-view-gadget. This may start the poll-process (like the menu-item ’select all’).

UNICONIFY

opens NetCall after ICONIFY.

16 NetCall 2.0c

5 User-interface

The mainwindow of NetCall displays two list-view-gadgets and one text-gadget.

All nodes, which should be called are displayed in the upper list-view-gadget. This are the nodes from your BOSS-entries (marked with an ’*’), the nodes from your CALL-entries and eventually nodes where you want to crash mails to or where you want to request files.

You can select those entries by clicking on them with the mouse or pressing one of the functionkeys (F1-F10). Selected nodes appear in the lower list-view-gadget.

Selected entries are displayed with all their lines (if they do have more than one line, of course). You can deselect individual entries by clicking on them with the mouse.

The entries in the list-view-gadgets have the following meaning:

  CRP *> 2:246/1045  (CDS Port #1)
  C : Crashmail ready for this node
  R : File-request waiting for this node
  P : This node waits to be polled (f.e. your boss)
  * : This entry is selected
  > : This entry is actually active, in other words: your mailer calls
      this node right now

The text-gadget displays actual information of the program.

18 NetCall 2.0c

6 Author information

You can always reach me as follows:

     Alexander Stirmlinger      Home: Weinbergweg 10
                                      D-74831 Gundelsheim
                           Phone/Fax: ++49 - 62 69 - 89 92
                               EMail: alex@picard.unterland.de
                             FidoNet: 2:246/1045.11

20 NetCall 2.0c

7 Acknowledgements

I’d like to thank the following people for their bug-reports, their many suggestions and their endless support during the beta-test(s):

Martin Biallas
  Rudi Brand
Frank Dürring
Klaus Gessner
 Stefan Hirth
Manfred Hoffmann
Peter Jachmann
Gerald Kleesiek
 Markus Mayer
  Sven Meyer
Alfred Schindler
Werner Senden
  Eric Wick

22 NetCall 2.0c

Appendix A How to register

If you want to use NetCall regularly, you have to register! If you register, you can use NetCall without any shareware-reminders. See also Section 1.4 [Restrictions of the unregistered version], Page 2.

Once registered, you are able to use all forthcoming versions of NetCall without sharewarereminders; your keyfile will be valid forever.

NetCall is shareware and costs 10,- DM (10.- US$). The following data is required for registration: your name, street, ZIP-code, city, telephone-

number and your email-address(es).

You may pay the share as follows:

1. Euro-cheque: DM (Deutsche Mark), payable to Alexander Stirmlinger.

2. Wire bank transfer: Please give NetCall as a comment. Please send me your address etc. via email or mail.

3. Cash: DM or US$ - at your own risk.

If you wish to get your keyfile PGP-encoded, send me your PGP-public first. If you want to get your keyfile by mail, please send me a blank disk and a self-addressed and

stamped envelope. See also Chapter 6 [Author info], Page 19.

Bank: Kreissparkasse Heilbronn Account : 190 3780 BLZ : 620 500 00

Thank you for supporting shareware!

24 NetCall 2.0c

Appendix B Support/Updates

You can always file-request or download the most recent version of NetCall at the CyberDyne System, Heilbronn, Germany (magic: NETCALL).

Registered users will receive release-notes as carbon-copies via email, if new NetCall-releases are available.

CyberDyne System (24h online):
Node 1: +49-7131-176169 * 2:246/1045 * 1200 - 16800 BPS
        V.32B V.42B ZYX
Node 2: +49-7131-161614 * 2:246/1046 * 1200 - 14400 BPS
        V.32B V.42B HST
Node 3: +49-7131-167890 * 2:246/1047 * 1200 - 28800 BPS
        V.32B V.42B V.FC
Node 4: +49-7131-950371 * 2:246/1048 * 64000 BPS BLASTER
        X.75 V.110 ISDN
Node 5: +49-7131-950372 * 2:246/1049 * 64000 BPS ELINK
        X.75 V.110 ISDN

26 NetCall 2.0c

Appendix C History

C.1 NetCall - Version 2.0

• First public release.

C.2 NetCall - Version 2.0a

• English documents included.

• New config-variable: MAXRETRIES.

• Minor ALLACTIVE-bug fixed.

• New shell-argument: TERMREQS. Supports automatical operation.

C.3 NetCall - Version 2.0b

• New config-variable: CALL.

• Minor bug fixed: no text was displayed in the text-gadget on startup under some conditions.

• New config-variable: BACKDROPWINDOW.

• New config-variable: STARTUPDELAY.

• Minor bug fixed: NetCall opened two requesters on exit (instead of exiting) under some conditions.

• Minor bug fixed: font adaption problem of the text-gadget solved.

• #?.REQ-file editor (’request file’) added.

• New config-variable: WINDOWFILEREQ.

• New config-variable: WINDOWFILEREQZOOM.

• New config-variable: AFTERSESSION.

• Now supports TrapDoor in resident mode, too.

• New config-variable: TRAPTELL.

• New shell-argument: TELL.

• New config-variable: SCREENTOFRONT.

• New config-variable: SUPRESSREQ.

• New config-variable: SAVELASTPOLLDATE.

• New menu-item: Rescan outbound.

• New menu-item: Add node entry.

• Expanded BOSS-config-variable.

• Changes in the OUTBOUND-directory will immediately become visible in the listview-gadgets of NetCall (notify-function).

• All entries in the listview-gadget can now be selected with the function-keys, too.

C.4 NetCall - Version 2.0c

• ’CONNECT’-requester in TELL-mode removed.

• The old config-parser had problems with spaces in the BOSS-entry ("BeforeSession" and "AfterSession") -> fixed.

• MAXRETRIES didn’t work correctly -> fixed.

• New config-variable: ADDNODE.

• Toolbox: NetCall features a window with user-defineable gadgets within. The functions of these gadgets can be defined by the user (start programms/scripts, send ARexx-commands etc.). Use the COMMAND-variable to configure the toolbox.

• New config-variable: WINDOWTOOLBOX.

• ARexx-interface (port: NETCALL).

• Installer-script will not overwrite the old NetCall.cfg now.

• If NetCall gets started a second time, a requester will appear (to bring the old screen of NetCall to the front).

• The list-view-gadget in the ’Select node...’-window (’Request file’) uses now 2-column output (looks nicer). (OS3.0 and above)

• Improved handling of very big fonts (GUI).

• NetCall now needs extacy.library >= V3.78.

• New config-variable: MAXRETRIESCOMMAND.

• New config-variable: REDIALDELAYENTRY.

• NodeListUpdate-script included in the archive.

Index 29

Index

A

ARexx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ARexx-interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Assumptions to the hardware . . . . . . . . . . . . . . . . . . . . . . . 3 Assumptions to the software . . . . . . . . . . . . . . . . . . . . . . . . 3 Author information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

C

Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

E

EMail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Execution from workbench . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

Features of NetCall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 FidoNet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

G

General informations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 GUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

H

History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

I

Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Installer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

K

Keyfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

M

MailManager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

N

NetCall.cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NetCall.polldates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 No warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

Paramenter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 5 PGP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

R

Restrictions of the unreg. version . . . . . . . . . . . . . . . . . . . 2

S

ShareWare. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Shell-arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Spot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

T

Tool-types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Toolbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 TrapDoor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 12 TrapDoor.cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

U

UMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 25 User-interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

W

Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Workbench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

30 NetCall 2.0c

Table of Contents

1 General informations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Features of NetCall. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Restrictions of the unregistered version . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Assumptions to the hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Assumptions to the software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Installer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Execution from a shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Execution from workbench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.1 The configuration-file NetCall.cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 The configuration of TrapDoor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.3 The usage of NetCall with Spot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 The usage of NetCall with UMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.5 The usage of NetCall with MailManager . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4 ARexx-interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5 User-interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6 Author information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Appendix A How to register . . . . . . . . . . . . . . . . . . . . . . . . . 23 Appendix B Support/Updates . . . . . . . . . . . . . . . . . . . . . . . 25 Appendix C History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

C.1 NetCall - Version 2.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 C.2 NetCall - Version 2.0a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 C.3 NetCall - Version 2.0b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 C.4 NetCall - Version 2.0c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29